12. Present value of annuities and annuity payments Aa Aa The present value of an annuity is the sum of the discounted value of all future cash flows. You have the opportunity to invest in several annuities. Which of the following 10-year annuities has the greatest present value (PV)? Assume that all annuities earn the same positive interest rate. O An annuity that pays $500 at the end of every six mońths O An annuity that pays $1,000 at the beginning of each year O An annuity that pays $1,000 at the end of each year O An annuity that pays $500 at the beginning of every six months An ordinary annuity selling at $15,513.53 today promises to make equal payments at the end of each year for the next four years (N). If the annuitys appropriate interest rate (1) remains at 5.00% during this time, the annual annuity payment (PMT) will be You just won the lottery. Congratulations! The jackpot is $60,000,000, paid in four equal annual payments. The first payment on the lottery jackpot will be made today. In present value terms, you really won assuming an annual interest rate of 5.00%. Flash Player MAC 31,0,0,108 engage Grade It Now
Answer #1

1.
An annuity that pays 1000 at the beginning of each year

2.
Using financial calculator:PV=-15513.53 N=4 I/Y=5% FV=0 CPT PMT=4374.999025
Using formula=15513.53*5%/(1-1/1.05^4)=4374.999025

3.
=60000000/(1-1/1.05^4)*(1-1/1.05)
=16114961.86306